A 1.00 m3 box is filled with 54.1 moles of air (MM = 0.0289 kg/mole). The rms speed of the molecules is 549 m/s. What is the pressure of the molecules? 


1
Expert's answer
2021-02-16T15:56:41-0500
"p=\frac {Nmv^2}{3V}"


"m=\nu\times MM=54.1\times0.0289=1.56 kg"

"N=\nu\times N_A=54.1\times 6.02\times10^{23}=3.26\times10^{25}"


"p=\frac {3.26\times 10^{25} \times 1.56 \times 549}{3\times 1}=9.3\times10^{27} Pa"